OpenStudy (anonymous):

write parabola in standard form: x^2-2x+8y+17=0

OpenStudy (anonymous):

So far i have attempted this problem but feel my work looks strange, I ended up with this as my answer: 8(y-(137/8)^2)=(x-1)^2

OpenStudy (anonymous):

standard form like \[(x-h)^2=4py\] that kind of standard form?

OpenStudy (anonymous):

yes @satellite73 that is correct!

OpenStudy (anonymous):

\[ x^2-2x+8y+17=0\] start with \[x^2-2x=-8y-17\] then complete the square on the left i have no idea how you got your answer!

OpenStudy (anonymous):

it would be (x-1)^2

OpenStudy (anonymous):

for the left side!

OpenStudy (anonymous):

yes and how about for the right?

OpenStudy (anonymous):

thats where i think i got stuck i wasnt sure what to do :|

OpenStudy (anonymous):

add 1!

OpenStudy (anonymous):

\[(x-1)^2=x^2-2x+1\] you added 1 to the left, you have to add 1 to the right

OpenStudy (anonymous):

thats right!

OpenStudy (anonymous):

so then,

OpenStudy (anonymous):

i would have -8y -16 .... -8(y+2)

OpenStudy (anonymous):

guess so yes

OpenStudy (anonymous):

wow not nearly as hard as all those other questions you usually ask!

OpenStudy (anonymous):

haha your always here to help me! this section is review and i haven't done this in years so its faded :|

OpenStudy (anonymous):

yeah that algebra gets you every time!

OpenStudy (anonymous):

sadly yes :| lol